3. Find all eigenvalues and eigenvectors for the following matrices R= [ { 1]
Answer #1

@ 101 A+ (3-47 D for eigenvalues IMI-A=0) =K? -0 7 (12)(+2)4=0 : 2120 5 2-4+4=0 7 10 7 130,0 for eigenvecfer do Consider (11-T (V) For det ca ayo (1791-01 for d= 1 CHI-A)y=0 = -7241 5 - 201-22=0,-) 00=-12 :V<= [i]), the